OpenStudy (anonymous):

Find the exact value of cos 315°.

OpenStudy (anonymous):

315 degrees is 45 degrees away from 0 degrees. Therefore sin(315)=cos(315). You also know that for any angle, x, sin%5E2%28x%29%2Bcos%5E2%28x%29=1 In this case, cos%5E2%28315%29%2Bcos%5E2%28315%29=1 2cos%5E2%28315%29=1 cos%5E2%28315%29=1%2F2 cos%28315%29=1%2Fsqrt%282%29 cos%28315%29=sqrt%282%29%2F2

OpenStudy (anonymous):

Crap the numbers dont work D:

OpenStudy (anonymous):

\[\frac{ \sqrt{2} }{ 2 }\] its C